You buy a 20-year bond with a coupon rate of 7.8% that has a yield to maturity of 9.9%. (Assume a face value of $1,000 and semiannual coupon payments.) Six months later, the yield to maturity is 10.9%. What is your return over the 6 months?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

AS NOTHING WAS MENTIONED, BA II PLUS USED TO SOLVE THE PROBLEM1 Ale $ w ? 1x ENG 23:05 06-02-2020 25 X fix BH151 2 146 BA BB BC BD BE BE BG BH BI 147 BA II PLUS set C/Y= P/Y = 148 149 150